the peremeter of a parallogram is 26 cm and the difference between its two adjecent side is 3cm taking the longer side as its base its area becomes 32 cm find the height of parallogrAM

Grade:8

1 Answers

Nitesh
6 years ago
Let base of the parallelogram x+3and it`s adjacent side is x Perimeter=2 (l+b)=》26=2 (x+x+3)=》x=5cm,x+3=8cmNow area of parallelogram =base×height=》32=8×height=》height=4 cmSo answer is 4 cm.